OpenStudy (anonymous):

Use the equation given below to find f '' (π/6). f (x) = csc(x)

OpenStudy (jdoe0001):

http://lh6.ggpht.com/sJzGWzl4LfQCY1O6P8PV7ccwY0Awj_vOw6IZzhj4ibHTqswhOIiykYc-Js7M9vlOeg=h900 get the derivative of csc(x) and then you'd end up with two factors so use the product rule for that

OpenStudy (anonymous):

okay i have got that with, csc^3x-cscxcot^2x, and plugged in the pi/6. however i cannot seem to get the right answer when i do it.

OpenStudy (loser66):

csc = 1/ sin csc' = (1/sin)' = -cos/sin^2 csc" = (-cos/sin^2)'= (-sin^3+ cos *(2sincos))/sin^4 = -1/sin +2cos^2/sin^3 replace pi/6 in to get f"= 10